OpenStudy (anonymous):

For an open-topped rectangular box, with a square base X by X cm and height h cm, find the dimensions giving the maximum volume, given that the surface area is 40cm6^2

OpenStudy (anonymous):

40cm^2

OpenStudy (mameko):

I am not 100% sure the answer, but here is what I think.  \[V = h * x ^{2}\] Volumn = hight * area of bottom \[S _{A} = 4hx + x ^{2}\] Surface area = 4 sides + bottom There are 2 variable we can use to take derivative . I choose x. so we have : \[S _{A}\prime = 4h + 4h _{x}\prime + 2x = \left( 40 \right)\prime = 0\] tidy up the equation we get: \[h _{x}\prime = \frac{ -x - 2h }{ 2x }\] Now we do the same thing to the volume : \[V _{x}\prime = h _{x}\prime x ^{2} + 2xh\] we get V max when the above equation at critical point. So we need to know when the above equation equal 0. Put hx ' in the Vx' equation and make it equal to 0. we get the relationship between x and h. 2h = x plug this back to the surface equation. you can get the actual value of x and h.